# ASK: Adaptive Self-improving Knowledge Framework for Audio Text Retrieval

### v1

@@ 4.1 Formulation of Knowledge Bases
Our framework’s first step is to construct multi-grained knowledge bases from a source dataset, $\mathcal{D}_{k}$. The choice of source is flexible; in our experiments, we explore three types to demonstrate versatility: 1) In-Domain + : the training set itself, 2) Out-of-Domain : WavCaps (mei2024wavcaps), and 3) Enriched In-Domain : training set re-annotated by Gemini (team2023gemini). From a chosen source, we construct two complementary bases.
@@
$\begin{split}K_{f}&=\{(u_{j}^{k},v_{j}^{k})\}_{j=1}^{N_{k}},\\ \text{where }u_{j}^{k}&=f_{\theta}(a_{j}^{k}),v_{j}^{k}=g_{\phi}(t_{j}^{k})\end{split}$ (6)
@@
$c_{m}^{u}=\text{MaxPool}(\{u_{j}^{k}\mid u_{j}^{k}\in\mathcal{C}_{m}^{u}\})$ (7)
@@
Figure 2: The proposed ASK framework. A multi-grained knowledge base ($K_{f},K_{c}$) is periodically updated to mitigate RDM. During training, knowledge is injected into samples ($u_{i}\to u^{\prime}_{i}$), and a cross-modal reliability weight ($\Psi$) is computed. A final loss is optimized using both an OT-realigned similarity matrix ($S^{*}$) and the reliability weight $\Psi$.
@@ 4.2 Multi-Grained Knowledge Injection
$\begin{split}u^{\prime}_{i,f}&=\rho u_{i}+(1-\rho)\bar{u}_{i}^{f},\\ \text{where }\bar{u}_{i}^{f}&=\frac{\sum_{(u_{j}^{k},v_{j}^{k})\in\mathcal{N}_{f}(u_{i})}u_{j}^{k}}{K}\end{split}$ (8)
@@
$\frac{\partial\mathcal{L}^{\prime}_{B}}{\partial u_{k}^{k}}=\sum_{i\in\mathcal{S}_{k}}\frac{\partial\mathcal{L}^{\prime}_{B}}{\partial u^{\prime}_{i,f}}\frac{\partial u^{\prime}_{i,f}}{\partial u_{k}^{k}}$ (9)
From Eq. 8, the second partial derivative is a non-zero constant $\frac{1-\rho}{K}$. Since the first derivative is also non-zero, the total gradient is non-zero. Consequently, the OBI, defined in Eq. 2, becomes strictly positive: $\text{OBI}(\mathcal{L}^{\prime}_{B})>0$. This quantitatively proves that our injection process breaks the GLB.
@@ 4.3 Adaptive Reliability Weighting
For a pair $(u_{i},v_{i})$, we start with the two sets of retrieved fine-grained neighbors: the audio-retrieved neighborhood for $u_{i}$, yielding a set of audio embeddings $\mathcal{U}_{r}=\{u_{l}^{k}\}_{l=1}^{K}$; and the text-retrieved neighborhood for $v_{i}$, yielding a set of audio-text pairs $\mathcal{N}_{f}(v_{i})=\{(u_{j}^{k^{\prime}},v_{j}^{k})\}_{j=1}^{K}$.
The process involves two main steps. First, we compute a consistency score $\bar{s}_{j}$ for each of the $K$ neighbors in $\mathcal{N}_{f}(v_{i})$. This score quantifies how well its audio partner $u_{j}^{k^{\prime}}$ aligns with the entire audio-retrieved neighborhood $\mathcal{U}_{r}$:
$\bar{s}_{j}=\frac{1}{K}\sum_{l=1}^{K}(u_{j}^{k^{\prime}})^{\top}u_{l}^{k}$ (10)
Second, these raw scores are normalized via a softmax function to form a probability distribution, representing a vector of reliability weights $\mathbf{w}_{f}=\{w_{j}\}_{j=1}^{K}$:
$w_{j}=\frac{\exp(\bar{s}_{j})}{\sum_{m=1}^{K}\exp(\bar{s}_{m})}$ (11)
These weights are then used to construct the final, reliability-aware knowledge potential, $\Psi_{i,f}^{T\to A}$. This potential is no longer a simple average, but a weighted sum of similarities, where each neighbor’s contribution is scaled by its cross-modal reliability:
$\Psi_{i,f}^{T\to A}=\sum_{j=1}^{K}w_{j}\cdot\exp(u_{i}^{\top}u_{j}^{k^{\prime}})$ (12)
A symmetric process is used to compute the text potential $\Psi_{i,f}^{A\to T}$, which measures the weighted alignment of the original text $v_{i}$ with the audio-retrieved text knowledge.
@@ 4.4 Dynamic Knowledge Refinement
As established in Section 3.3, training with a static knowledge base is fundamentally unstable due to the Representation Drift Mismatch (RDM), which increases the risk of gradient misalignment over time. To counteract this, we introduce a dynamic knowledge refinement mechanism.
This mechanism periodically reconstructs the knowledge bases $K_{f}$ and $K_{c}$ using the current model encoders. The process is governed by an update period hyperparameter $\mathcal{T}$, which defines the number of training epochs between each refinement.
The theoretical justification for this is its direct impact on the RDM. At each update step $t$, the refinement process effectively resets the knowledge base timestamp $t_{k}$ to $t$. As a result, the ideal and actual neighborhood distributions (Eq. 3) become identical, $P_{\text{ideal}}\equiv P_{\text{actual}}$. Consequently, the RDM, as defined in Eq. 4, is reset to its optimal value:
$\text{RDM}(t,t)=\mathbb{E}[D_{KL}(P_{\text{ideal}}||P_{\text{ideal}})]=0$ (13)
By periodically resetting the RDM, our mechanism also resets the upper bound on gradient deviation (Eq. 5), thus ensuring a robust and stable training process where the knowledge base co-evolves with the model.
@@ 4.5 Unified Optimization Objective
The process begins with the knowledge-enhanced embeddings from Section 4.2. For a mini-batch, we compute a fine-grained similarity matrix $\mathbf{S}_{f}$ and a coarse-grained one $\mathbf{S}_{c}$. Since the audio and text knowledge are retrieved independently, the distributions of their nearest neighbors within the batch may differ. To reconcile this potential discrepancy and find a globally optimal batch-level matching, we employ Optimal Transport (OT) (cuturi2013sinkhorn) to learn an optimal transport plan $\mathbf{Q}^{*}$ (the full formulation is detailed in Appendix C). This plan is then used to produce the realigned similarity matrices $\mathbf{S}^{*}_{f}$ and $\mathbf{S}^{*}_{c}$:
$\mathbf{S}^{*}_{f}=\big((1-\beta)\mathbf{I}+\beta\,\mathbf{Q}^{*}\big)\,\mathbf{S}_{f}$ (14)
@@
$\begin{split}\mathcal{F}_{f}^{T\to A}&=\frac{1}{|B|}\sum_{i=1}^{|B|}-\log\Psi_{i,f}^{T\to A}\\ \mathcal{F}_{c}^{T\to A}&=\frac{1}{|B|}\sum_{i=1}^{|B|}-\log\Psi_{i,c}^{T\to A}\end{split}$ (16)
@@
$\mathcal{L}_{\text{ASK}}=\frac{1}{2}(\mathcal{L}^{*}_{T\to A}+\mathcal{L}^{*}_{A\to T})$ (18)
This composite objective ensures the model learns from multi-grained knowledge that is both globally aligned at the batch level and weighted by its cross-modal reliability. Furthermore, we provide a theoretical proof in Appendix B that demonstrates the convergence properties of our ASK framework.
@@ 5 Experiments
Table 1: Results for Audio-Text-Retrieval on AudioCaps and Clotho under the global interaction strategy. The symbols +, , and denote the use of knowledge from WavCaps, the Gemini-annotated training set, and the original training set, respectively.
@@
Method A2T T2A A2T T2A
@@
Architecture: ResNet-38 + BERT
mei2022metric $36.3_{\pm 0.5}$ $68.6_{\pm 0.3}$ $81.5_{\pm 0.2}$ $32.2_{\pm 0.4}$ $68.2_{\pm 0.1}$ $81.2_{\pm 0.2}$ $16.3_{\pm 0.4}$ $39.1_{\pm 0.3}$ $51.5_{\pm 0.6}$ $14.2_{\pm 0.4}$ $37.3_{\pm 0.2}$ $49.9_{\pm 0.3}$
@@
Architecture: CED-Base + SONAR-TE
yan2024bridging $39.6_{\pm 0.2}$ $69.8_{\pm 0.3}$ $81.7_{\pm 0.6}$ $31.9_{\pm 0.3}$ $69.2_{\pm 0.5}$ $82.8_{\pm 0.9}$ $18.0_{\pm 0.2}$ $39.5_{\pm 0.7}$ $53.0_{\pm 0.6}$ $14.9_{\pm 0.3}$ $39.9_{\pm 0.6}$ $53.1_{\pm 0.7}$
@@ 5.1 Experimental Setup
We evaluate our method on two standard benchmarks: AudioCaps (kim2019audiocaps) and Clotho (drossos2020clotho). Following prior work (mei2022metric; xie2024gpa; yan2024bridging), we report audio-to-text (A2T) and text-to-audio (T2A) retrieval performance using Recall at K (R@K, for K=1, 5, 10).
@@
To validate the model-agnostic nature of ASK, we integrate it into two types of baselines. 1) Global Interaction: We use a PANNs-based ResNet-38 (kong2020panns) + BERT (devlin2019bert) pair (mei2022metric), and a ViT-based CED-Base (dinkel2024ced) + SONAR-TE (duquenne2023sonar) pair, following ML-CLAP (yan2024bridging) but trained only on English. 2) Local Interaction: We adapt the GPA (xie2024gpa) setup, using its ResNet-38 + BERT architecture but removing the Sinkhorn inference module to form a strong baseline, and set the same maximum number of tokens for the entire dataset.
@@
All models are trained with the Adam optimizer (adam2014method). For the ResNet-BERT architecture, we train for 50 epochs on AudioCaps (batch size 32) and Clotho (batch size 24) with an initial learning rate of $5\times 10^{-5}$, which is decayed by a factor of 10 every 20 epochs. The CED-SONAR models are trained for 10 epochs with a decay every 4 epochs. We use the Faiss library (douze2025faiss) for efficient neighbor search. Unless specified otherwise, the hyperparameters for our ASK framework are set as follows: we retrieve $K=10$ neighbors, with a coarse-grained prototype set of size $N_{c}=512$. The knowledge injection ratio is $\rho=0.2$, and the OT-alignment factor is $\beta=0.2$. The reliability modulation weights are $\lambda_{f}=0.2$ and $\lambda_{c}=0.3$. The knowledge base is dynamically refined every $\mathcal{T}=15$ epochs. All experiments were conducted on 2 NVIDIA A100 and 8 RTX 4090 GPUs.

### v2

@@ 4.1 Formulation of Knowledge Bases
Our framework’s first step is to construct multi-grained knowledge bases from a source dataset, $\mathcal{D}_{k}$. The choice of source is flexible; in our experiments, we explore three types to demonstrate versatility: 1) In-Domain + : the training set itself, 2) Out-of-Domain : WavCaps (Mei et al., 2024), and 3) Enriched In-Domain : training set re-annotated by Gemini 2.5 (Comanici et al., 2025). From a chosen source, we construct two complementary bases.
@@
$\begin{split}K_{f}&=\{(u_{j}^{k},v_{j}^{k})\}_{j=1}^{N_{k}},\end{split}$ (7)
where $u_{j}^{k}=f_{\theta}(a_{j}^{k}),v_{j}^{k}=g_{\phi}(t_{j}^{k})$.
@@
$c_{m}^{u}=\text{MaxPooling}(\{u_{j}^{k}\mid u_{j}^{k}\in\mathcal{C}_{m}^{u}\}).$ (8)
@@ 4.2 Multi-Grained Knowledge Injection
$\begin{split}u^{\prime}_{i,f}&=(1-\rho)u_{i}+\rho\bar{u}_{i}^{f},\\ \bar{u}_{i}^{f}&=\frac{\sum_{(u_{j}^{k},v_{j}^{k})\in\mathcal{N}_{f}(u_{i})}u_{j}^{k}}{K},\end{split}$ (9)
@@
$\frac{\partial\mathcal{L}^{\prime}_{B}}{\partial u_{k}^{k}}=\sum_{i\in\mathcal{S}_{k}}\frac{\partial\mathcal{L}^{\prime}_{B}}{\partial u^{\prime}_{i,f}}\frac{\partial u^{\prime}_{i,f}}{\partial u_{k}^{k}}.$ (10)
From Eq. 9, the second partial derivative is a non-zero constant $\frac{\rho}{K}$. Given that one of its partial derivatives is non-zero, the total gradient is therefore non-zero. Consequently, the OBI, defined in Eq. 2, becomes strictly positive. This quantitatively proves that our injection process breaks the GLB.
@@ 4.3 Adaptive Reliability Weighting
For each pair $(u_{i},v_{i})$, we consider two fine-grained neighborhoods: the audio-retrieved audio set $\mathcal{U}_{r}=\{u_{l}^{k}\}_{l=1}^{K}$ and the text-retrieved audio–text set $\mathcal{N}_{f}(v_{i})=\{(u_{j}^{k^{\prime}},v_{j}^{k})\}_{j=1}^{K}$. We first assign each neighbor in $\mathcal{N}_{f}(v_{i})$ a consistency score $\bar{s}_{j}$, defined as its average similarity to the audio-retrieved neighborhood:
$\bar{s}_{j}=\frac{1}{K}\sum_{l=1}^{K}(u_{j}^{k^{\prime}})^{\top}u_{l}^{k}.$ (11)
These scores are subsequently normalized via a softmax function to yield the reliability weights $\mathbf{w}_{f}=\{w_{j}\}_{j=1}^{K}$:
$w_{j}=\frac{\exp(\bar{s}_{j})}{\sum_{m=1}^{K}\exp(\bar{s}_{m})}.$ (12)
The reliability-aware knowledge potential is then computed as the weighted similarity between $u_{i}$ and the audio components of $\mathcal{N}_{f}(v_{i})$:
$\Psi_{i,f}^{T\to A}=\sum_{j=1}^{K}w_{j}\cdot\exp(u_{i}^{\top}u_{j}^{k^{\prime}}).$ (13)
A symmetric construction produces the text-side potential $\Psi_{i,f}^{A\to T}$, based on the audio-retrieved text neighborhood.
@@ 4.4 Dynamic Knowledge Refinement
As shown in Section 3.3, a static knowledge base leads to Representation Drift Mismatch (RDM), which induces increasing gradient misalignment during training. To mitigate this, we employ a dynamic refinement mechanism that periodically reconstructs the knowledge bases $K_{f}$ and $K_{c}$ using the current encoders. The update period $\mathcal{T}$ specifies the number of epochs between successive reconstructions.
This procedure directly controls the RDM. At each update step $t$, refinement sets the knowledge-base timestamp to $t_{k}=t$, making the ideal and actual neighborhood distributions identical, $P_{\text{ideal}}\equiv P_{\text{actual}}$. Thus, the RDM (Eq. 4) is reset to its minimum value:
$\text{RDM}(t,t)=\mathbb{E}[D_{KL}(P_{\text{ideal}}\,\|\,P_{\text{ideal}})]=0.$ (14)
By periodically driving the RDM to zero, the mechanism also resets the upper bound on gradient deviation (Eq. 6), ensuring stable optimization and enabling the knowledge base to co-evolve with the model.
@@ 4.5 Unified Optimization Objective
The process begins with the knowledge-enhanced embeddings from Section 4.2. For a mini-batch, we compute a fine-grained similarity matrix $\mathbf{S}_{f}$ and a coarse-grained one $\mathbf{S}_{c}$. Since the audio and text knowledge are retrieved independently, the distributions of their nearest neighbors within the batch may differ. To reconcile this potential discrepancy and find a globally optimal batch-level matching, we employ Optimal Transport (OT) (Cuturi, 2013) to learn an optimal transport plan $\mathbf{Q}^{*}$. This plan is then used to produce the realigned similarity matrices $\mathbf{S}^{*}_{f}$ and $\mathbf{S}^{*}_{c}$:
$\mathbf{S}^{*}_{f}=\big((1-\beta)\mathbf{I}+\beta\,\mathbf{Q}^{*}\big)\,\mathbf{S}_{f}.$ (15)
@@
$\begin{split}\mathcal{F}_{f}^{T\to A}&=\frac{1}{|B|}\sum_{i=1}^{|B|}-\log\Psi_{i,f}^{T\to A},\\ \mathcal{F}_{c}^{T\to A}&=\frac{1}{|B|}\sum_{i=1}^{|B|}-\log\Psi_{i,c}^{T\to A}.\end{split}$ (17)
@@
$\mathcal{L}^{*}_{T\to A}=(1+\lambda_{f}\mathcal{F}_{f}^{T\to A}+\lambda_{c}\mathcal{F}_{c}^{T\to A})\cdot\mathcal{L}_{T\to A},$ (18)
@@
$\mathcal{L}_{\text{ASK}}=\frac{1}{2}(\mathcal{L}^{*}_{T\to A}+\mathcal{L}^{*}_{A\to T}).$ (19)
This composite objective ensures the model learns from multi-grained knowledge that is both globally aligned at the batch level and weighted by its cross-modal reliability.
4.6 Theoretical Analysis
We provide a theoretical justification for ASK, framing the training as an alternating optimization to maximize the log-likelihood of observed audio-text pairs $x_{i}=(a_{i},t_{i})$.
Probabilistic Formulation via ELBO
Our goal is to maximize the log-likelihood $\mathcal{L}(\theta)=\sum_{i}\log p(x_{i};\theta)$. By introducing latent variables $z_{i}=(z_{i,f},z_{i,c})$ representing the optimal knowledge, and an auxiliary distribution $Q(z_{i})$, we apply Jensen’s Inequality to derive the Evidence Lower Bound (ELBO), $\mathcal{F}(Q,\theta)$:
$\mathcal{L}(\theta)$ $=\sum_{i}\log\sum_{z_{i}}\frac{p(x_{i},z_{i};\theta)}{Q(z_{i})}Q(z_{i})$ (20)
$\geq\sum_{i}\mathbb{E}_{Q(z_{i})}[\log p(x_{i},z_{i};\theta)]+H(Q)\triangleq\mathcal{F}(Q,\theta),$
where $H(Q)$ is the entropy and maximizing $\mathcal{L}(\theta)$ is achieved by iteratively maximizing $\mathcal{F}(Q,\theta)$.
Alternating Optimization.
Let $\theta_{t}$ be the parameters at iteration $t$. The process alternates between two stages:
Stage 1: Auxiliary Distribution Update. Fixing $\theta_{t}$, we approximate the optimal $Q_{t}(z_{i})$ using the retrieved neighbors. We define the probability mass of $Q_{t}$ over a neighbor $z_{j}$ directly via our reliability weights (Eq. 12): $Q_{t,f}(z_{j}):=w_{j,f}(\theta_{t})$ and $Q_{t,c}(z_{j}):=w_{j,c}(\theta_{t})$.
Stage 2: Model Parameter Update. Fixing $Q_{t}$, we maximize the expectation $\mathbb{E}_{Q_{t}}[\log p(x_{i},z_{i};\theta)]$. We model the joint log-probability as the negative sum of the alignment loss and the reliability potential: $\log p(x_{i},z_{i};\theta)\propto-(\mathcal{L}_{OT}(\theta)+\log\Psi_{i}(\theta))$. Substituting this into the ELBO, the maximization objective becomes minimizing the negative expectation:
$\min_{\theta}\mathcal{L}_{m}\approx\sum_{i}\left(\mathbb{E}_{Q_{t,f}}[\mathcal{L}_{OT,f}+\log\Psi_{i,f}]+\mathbb{E}_{Q_{t,c}}[\mathcal{L}_{OT,c}+\log\Psi_{i,c}]\right).$ (21)
This objective $\mathcal{L}_{m}$ mathematically aligns with our modulated loss $\mathcal{L}^{*}$ (Eq. 18), where the reliability term $\mathcal{F}=-\log\Psi$ acts as a regularizer.
Convergence.
Since each step monotonically increases the ELBO $\mathcal{F}(Q,\theta)$, and $\mathcal{L}_{\text{ASK}}$ is bounded below, the Monotone Convergence Theorem guarantees that the sequence of loss values converges to a stationary point.
@@ 5 Experiments
Table 1: Results for Audio-Text-Retrieval on AudioCaps and Clotho under the global interaction strategy. The symbols , , and + denote the use of knowledge from WavCaps, the Gemini-annotated training set and the original training set respectively.
@@
Method Audio-to-Text Text-to-Audio Audio-to-Text Text-to-Audio
@@
[…]
ML-ACT Mei et al. (2022) $36.3_{\pm 0.5}$ $68.6_{\pm 0.3}$ $81.5_{\pm 0.2}$ $32.2_{\pm 0.4}$ $68.2_{\pm 0.1}$ $81.2_{\pm 0.2}$ $16.3_{\pm 0.4}$ $39.1_{\pm 0.3}$ $51.5_{\pm 0.6}$ $14.2_{\pm 0.4}$ $37.3_{\pm 0.2}$ $49.9_{\pm 0.3}$
BLAT Xu et al. (2023) $38.2_{\pm 0.2}$ $70.4_{\pm 0.3}$ $82.1_{\pm 0.2}$ $32.9_{\pm 0.3}$ $68.9_{\pm 0.1}$ $81.8_{\pm 0.2}$ $16.8_{\pm 0.2}$ $39.6_{\pm 0.3}$ $52.1_{\pm 0.3}$ $14.1_{\pm 0.2}$ $37.6_{\pm 0.2}$ $50.2_{\pm 0.1}$
Auto-ACD Sun et al. (2024) $40.8_{\pm 0.2}$ $71.3_{\pm 0.4}$ $83.3_{\pm 0.2}$ $33.2_{\pm 0.3}$ $68.7_{\pm 0.2}$ $82.1_{\pm 0.2}$ $17.1_{\pm 0.2}$ $39.3_{\pm 0.2}$ $53.2_{\pm 0.3}$ $14.4_{\pm 0.2}$ $37.5_{\pm 0.2}$ $50.1_{\pm 0.2}$
@@
[…]
ML-CLAP Yan et al. (2024) $39.6_{\pm 0.2}$ $69.8_{\pm 0.3}$ $81.7_{\pm 0.6}$ $31.9_{\pm 0.3}$ $69.2_{\pm 0.5}$ $82.8_{\pm 0.9}$ $18.0_{\pm 0.2}$ $39.5_{\pm 0.7}$ $53.0_{\pm 0.6}$ $14.9_{\pm 0.3}$ $39.9_{\pm 0.6}$ $53.1_{\pm 0.7}$
GLAP Dinkel and others (2025) $41.2_{\pm 0.1}$ $72.4_{\pm 0.3}$ $83.7_{\pm 0.2}$ $33.3_{\pm 0.3}$ $68.9_{\pm 0.1}$ $81.8_{\pm 0.4}$ $18.4_{\pm 0.2}$ $40.6_{\pm 0.3}$ $54.1_{\pm 0.3}$ $15.1_{\pm 0.2}$ $40.2_{\pm 0.2}$ $54.2_{\pm 0.1}$
@@ 5.1 Experimental Setup
We evaluate our method on two standard benchmarks: AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020). Following prior work (Mei et al., 2022; Xie et al., 2024; Yan et al., 2024), we report audio-to-text (A2T) and text-to-audio (T2A) retrieval performance using Recall at K (R@K, for K=1, 5, 10).
@@
To validate ASK’s model-agnosticism, we evaluate it across two interaction paradigms. 1) Global Interaction: We build upon the ML-ACT baseline (Mei et al., 2022) (ResNet-38 (Kong et al., 2020) + BERT (Devlin et al., 2019)), comparing against BLAT (Xu et al., 2023) and Auto-ACD (Sun et al., 2024). We also adapt the English-only ML-CLAP setup (Yan et al., 2024) (CED-Base (Dinkel et al., 2024) + SONAR-TE (Duquenne et al., 2023)), comparing it with GLAP (Dinkel and others, 2025). 2) Local Interaction: We follow the setups of GPA (Xie et al., 2024) and FLAM (Wu et al., ), and adopt the same maximum number of tokens for the dataset.
@@
All models are trained with the Adam optimizer (Kingma and Ba, 2014). The ResNet-BERT architecture is trained for 50 epochs on AudioCaps (batch size 32) and Clotho (batch size 24), with an initial learning rate of $5\times 10^{-5}$, which is decayed by a factor of 10 every 20 epochs. The CED-SONAR models are trained for 10 epochs with a decay step applied every 4 epochs. We use the Faiss library (Douze et al., 2025) for efficient neighbor search. Unless specified otherwise, the hyperparameters for our ASK framework are set as follows: we retrieve $K=10$ neighbors, with a coarse-grained prototype set of size $N_{c}=512$. The knowledge injection ratio is $\rho=0.2$, and the OT-realignment factor is $\beta=0.2$. Empirical analyses confirm our framework’s robustness across $K\in[5,15]$ and $\rho\in[0.15,0.25]$. The reliability modulation weights are $\lambda_{f}=0.2$ and $\lambda_{c}=0.3$. The knowledge base is dynamically refined every $\mathcal{T}=15$ epochs. All experiments were conducted on 2 NVIDIA A100 and 8 RTX 4090 GPUs.